OpenStudy (anonymous):

Simplify completely. √18y^10

OpenStudy (anonymous):

√18y^10 = √18*√y^10 =3√2*y^(10/2) =(3√2)*y^5

OpenStudy (anonymous):

So what your saying is 3y^5√2 ???

OpenStudy (anonymous):

@nukka

OpenStudy (anonymous):

Yeah that's what it will be expanded.

OpenStudy (anonymous):

As long as the question is √18y^10 and not √(18y)^10

OpenStudy (anonymous):

ok